OpenStudy (anonymous):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks. Number of pages 50 100 150 200 Cost 1.75 2.55 3.35 4.15

OpenStudy (kamibug):

Okay. What's your question? :)

OpenStudy (anonymous):

Oooh lol forgot it >.<

OpenStudy (anonymous):

Which equation gives the cost, y, in terms of the number of pages, x?

OpenStudy (kamibug):

Are there options? :)

OpenStudy (anonymous):

A. y = 0.016x + 0.95 B. y = 0.016x + 1.75 C. y = 0.016x + 48.25 D. y = 52.6x + 0.95

OpenStudy (anonymous):

@kamibug?

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

i neeed help :(

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

@DanJS